OpenStudy (anonymous):

Jack usually mows his lawn in 6 hrs. Marilyn can mow the same yard in 7 hrs. How much time would it take for them to mow the lawn together?

OpenStudy (dumbcow):

6x + 7x = 42 13x = 42 x = 42/13 = 3.23 hours explanation: find hourly rate of each person Marilyn works at rate of 1/7 yard per hour Jack at a rate of 1/6 yard per hour 1/7 x + 1/6 x = 1 multiply equation by 42 (LCM of 6,7) 6x + 7x = 42